Pluto Travel

Contribution Margin Income Statement

Three Months Ended March 31, 2016

Net Sales Revenue

$318,500

Variable Costs

111,475

Contribution Margin

207,025

Fixed Costs

170,000

Operating Income

$37,025

For its top​ managers, Pluto Travel formats its income statement as​ follows: Pluto's relevant range is between sales of $256,000 and $362,000.

1.

Calculate the contribution margin ratio.

2.

Prepare two contribution margin income​ statements: one at the $256,000 sales level and one at the $362,000 sales level.

​(Hint: The proportion of each sales dollar that goes toward variable costs is constant within the relevant​ range.)
Answer #1

Page No ① Answer: I The Contribution Margin Ratio - contribution Hangin Ratio : cion Margin Ratio . Contribution dl 207025 31Pageno =89600 · Sales = $ 362,000 Variable Cost - 362000-(362000 x 654) = 362000 - 205 300 : = 126700 sales 256000 362000 Var